Four pennies of diameter d are placed with their centers on the four corners of a d x d square. the percentage of the area of the square covered by the pennies is
A 60 to 70 b 70 to 80 c 80 to 90 d 90 to 100 e none of these

Answer by richwmiller(9144)

we have 4*1/4 =1 penny
area of a penny pi*(d/2)^2=3.14*d^2/4=3.14/4*d^2
area of the square is d^2
3.14/4 *100 is the percentage of the square covered by the pennies
3.14/4 *100=78.5
so b) 70-80